Ledio and Isabel Greene

5.Ledio and Isabel Greene have the following assets, liabilities, income, and expenses. Use this information to answer the questions that follow:

•Yearly income: $99,000

•Value of home: $190,000

•Value of cars: $32,000

•Monetary Assets: $30,000

•Mortgage on house: $150,000 with twenty-four years remaining on the loans

•Auto loans: $14,000 with three years remaining on the loans

•Student loans: $92,000 with five years remaining on the loans

•Mortgage payment: $1,200

•Utilities: $500

•Car payments: $550

•Groceries: $900

•Taxes: $3,000

•Discretionary expenses: $1,000

a.What is their current net worth situation?

b.What is their current discretionary cash flow position?

c.What is their current ratio?

d.What is their debt ratio?

e.What is their savings ratio (assume that any discretionary cash flow is used for savings)?

f.Ledio and Isabel are considering purchasing a new car. The car will cost $30,000. They plan to borrow the entire amount of the car purchase. By how much will their net worth increase or decrease?
Answer #1

Part (a)

Assets Linkage $
Home A 190,000
Cars B 32,000
Monetary Assets C 30,000
Total Assets D = A + B + C 252,000
Liabilities
Mortgage on house E        150,000
Auto loan F           14,000
Student loan G           92,000
Total Liabilities H = E + F + G        256,000
Part (a) Current networth situation D - H           (4,000)

Part (b)

Monthly Income A 8,250
Mortgage payment B 1,200
Utilities C 500
Car payments D 550
Groceries E 900
Taxes F 3,000
Part (b) Current discretionary cash flow position G = A - Sum of B to F 2,100

Part (c)

Annual expenses = 12 x monthly expenses = 12 x (1,200 + 500 + 550 + 900 + 3,000) = 73,800

Current Ratio = Monetary assets / Annual expenses = 30,000 / 73,800 = 0.4065

Part (d)

Debt ratio = total liabilities / total assets =  256,000 / 252,000 = 1.0159

Part (e)

Savings ratio = Discretionary position / income = 2,100 / 8,250 = 0.2545

Part (f)

Increase in net worth = Asset - liability = 30,000 - 30,000 = 0

Hence, the net worth will not change. It will remain the same.